OpenStudy (anonymous):

***WILL FAN AND MEDAL*** Gina recently got a %10 raise on her salary. If Gina used to make $25 an hour, what is her new salary? I just need to know the formula.

OpenStudy (anonymous):

@jabez177

OpenStudy (abdullahm):

Two ways to do this: Multiply 25 by 10% and then add 25 to it. Or multiply 25 by 110% since we are increasing it by 10% Remember 10% = 10/100 110% = 110/100